# From Code to Correctness: Closing the Last Mile of Code Generation with Hierarchical Debugging

### v3

@@ Abstract.
copyright: none
@@ 1. Introduction
Large language models (LLMs) such as GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023), and DeepSeek-Coder (AI, 2024) have made significant advances in AI-assisted coding tasks (Chen et al., 2021; Li et al., 2022; Just et al., 2014). Trained on vast corpora of text and code, LLMs can understand and generate code snippets for various programming tasks, ranging from simple data structures to complex algorithmic problems (Li et al., 2022). These models have demonstrated proficiency in tasks such as code completion, bug detection, and even tackling competitive programming challenges (Chen et al., 2021; Zhang et al., 2024b; Wen et al., 2024).
While the code generated by large language models generally meets the requirements, it often contains critical errors that require human intervention to pass tests (Liu et al., 2023a; Dou et al., 2024). This has gradually led to a new development paradigm: large models generate the code, while humans fix it. Therefore, the “last mile”, as well as the most crucial step, of code generation is how to efficiently repair the code generated by large models.
Numerous efforts have been made to debug LLM-generated code. The most popular way is to reuse the LLM generator to debug the generated code with the feedback from test case execution (Chen et al., 2023b; Zhong et al., 2024; Hu et al., 2024). While these methods increase the pass rates, they treat the erroneous program as a holistic set of statements (i.e., as a single unit without considering the internal structure or modularity) (Chen et al., 2023b; Shinn et al., 2023; Zhong et al., 2024; Ding et al., 2024) regardless of the varying types and levels of failures. Failures of test cases arise from different levels of factors, from low-level syntactic errors to high-level algorithmic flaws. A holistic treatment overlooks the internal structure of the code and limits the effectiveness of the debugging systems, especially when dealing with complex programs that need debugging across different modules (Zeller, 2009; Tian et al., 2024).
@@
Extensive experiments across multiple models and benchmarks demonstrate that MGDebugger significantly outperforms existing debugging methods. Using open-source models like DeepSeek-Coder-V2-Lite, CodeQwen1.5, and Codestral, MGDebugger elevates accuracy from 75.6% to 94.5% on HumanEval (Chen et al., 2021) and achieves a remarkable 97.6% repair success rate on HumanEvalFix (Muennighoff et al., 2023). The approach also generalizes well to state-of-the-art proprietary models, improving GPT-4o and Claude 3.5 Sonnet’s performance by 6.1% on HumanEval, reaching accuracies of 96.3% and 95.7% respectively.
@@
While MGDebugger is primarily designed to fix LLM-generated code and close the last mile of code generation, we also evaluate its generalizability to real-world software defects. When applied to the Defects4J dataset (Just et al., 2014), MGDebugger successfully repairs 129 bugs in Defects4J v1.2 and 64 bugs in Defects4J v2.0, outperforming specialized program repair tools. This demonstrates that the hierarchical debugging principles with test case generation and LLM-simulated execution of MGDebugger extend effectively beyond LLM-generated code to human-written software with complex project structures.
@@ 2.2. Hierarchical Code Decomposition
Modularizing and decomposing complex code into smaller helper subfunctions has been proven to be helpful especially for large functions that are difficult to understand (Jain et al., 2023; Zelikman et al., 2023). To enable hierarchical debugging, we need to transform the input code into a tree-like structure of subfunctions.
Specifically, given an LLM-generated function $f$, we decompose it into a hierarchical structure of subfunctions denoted as $(f_{1},...,f_{n})$. We formally define a ”subfunction” as a contiguous sequence of statements $S=[s_{1},s_{2},...,s_{k}]$ within function $f$, such that $S$ represents a logically coherent and minimal unit of functionality that can be abstracted as a function $f_{i}(x)$ with well-defined inputs and outputs. These subfunctions can be organized as a tree $f_{\text{root}}=\text{TREE}(f_{\text{root}},\text{CHILD}(f_{\text{root}}))$, where $f_{\text{root}}$ represents the main function and $\text{CHILD}(f)$ denotes the set of subfunctions directly called by $f$. We leverage an LLM for the decomposition, adhering to three principles: (1) each subfunction represents the minimal reusable unit of code with a specific purpose, (2) higher-level functions call lower-level functions to achieve complex functionality, and (3) the overall structure facilitates isolated testing and debugging. As illustrated in Figure 1, the resulting tree-like structure allows us to isolate logical units of the code, enabling more focused debugging efforts across different levels of granularity (Woodfield et al., 1981; Isazadeh et al., 2017). The prompt template for hierarchical decomposition is shown below:
@@ 2.3. Generating Test Cases for Subfunctions
Having obtained the hierarchy of subfunctions, we aim to verify the correctness of each subfunction. For this purpose, we generate test cases for each subfunction leveraging automatic unit test generation techniques (Wang et al., 2021; Schäfer et al., 2024; Liu et al., 2023b). For each subfunction $f_{i}\in f_{\text{root}}$, we generate a set of test cases $\mathcal{T}_{i}$. Following the problem settings from Chen et al. (2023b) and Zhong et al. (2024), we assume that the public test cases for the main function $\mathcal{T}_{\text{pub}}$ have been provided, which is common in most code generation benchmarks (Chen et al., 2021; Hendrycks et al., 2021; Muennighoff et al., 2023) [^2] Otherwise, we can use LLM-generated test cases instead.. We can leverage these test cases to derive a set of corresponding test cases for each subfunction.
@@ 2.4. Debugging Subfunctions with LLM-simulated Execution
One straightforward way to implement this process is to use an external Python executor to monitor runtime variable values (Zhong et al., 2024). However, when debugging high-level functions, tracking variable values within lower-level subfunctions is often unnecessary, as their correctness is ensured by the bottom-up debugging methodology. Furthermore, directly collecting all execution traces from the external debugger can add unnecessary overhead and complexity to the process.
Inspired by the methodology in Li et al. (2023), we propose an LLM-simulated code executor, which prompts LLM to act as a code interpreter and track the code execution. We choose LLM-simulated execution over real execution for several key reasons (La Malfa et al., 2024; Ni et al., 2024; Zheng et al., 2024): (1) Many code snippets require specific environments, hardware, or dependencies that are difficult to set up or may not be available (e.g., embedded systems, special libraries), (2) Some code cannot be executed due to compilation errors or security concerns, necessitating sandboxing or virtual execution. As shown in Figure 2, we request the LLM to simulate the execution process, reasoning about key variables and their states at each step, and thoroughly analyzing the failed test cases. This eliminates the need for an external debugger, offering a more flexible and efficient debugging solution. In addition, the LLM can accurately identify where errors occur and grasp their surrounding context. The prompt template for debugging subfunction is shown below:
@@
12:  if pass ($\mathcal{R}_{f}$, $\mathcal{T}_{f}$then
@@ 3.1. Setup
Models We select three state-of-the-art open-source LLMs ranging from 7B to 22B parameters as backbones for code generation and debugging in the experiments: CodeQwen1.5 (7B) (Team, 2023), DeepSeek-Coder-V2-Lite (16B) (AI, 2024), and Codestral (22B) (Mistral AI team, 2024). To validate the generalizability of MGDebugger, stronger general-prupose models like GPT-4o (OpenAI, 2023), Claude 3.5 Sonnet (Anthropic, 2024), and LLaMA 3.1 70B (Touvron et al., 2023) are also included in the Experiments.
@@
Datasets We conduct experiments on three widely used benchmarks for LLM-based code generation and debugging. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are two widely used benchmarks for evaluating code generation systems with 164 and 500 problems, respectively. The HumanEvalFix dataset (Muennighoff et al., 2023) consists of 164 buggy functions with six different bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse.
@@
Metrics We adopt two metrics to evaluate our method: 1) Accuracy (Chen et al., 2023b; Zhong et al., 2024), which measures the overall proportion of correct code samples among all generated code samples after debugging. A code is correct iff it passes all private test cases assigned to it. 2) Repair Success Rate (RSR) (Yasunaga and Liang, 2021), which refers to the proportion of fixed code samples to the total number of buggy code samples. The former metric indicates to which extent MGDebugger can improve the code generation systems, and the latter metric directly reflects the effectiveness of MGDebugger in fixing bugs from LLM-generated code.
@@
Baselines We compare MGDebugger with eight state-of-the-art methods for debugging LLM-generated code. 0) No-Debugging baseline refers to directly using the LLM to generate code without applying any debugging method, following (Chen et al., 2023b; Zhong et al., 2024). This baseline creates the initial set of seed programs, the subset of buggy programs then become targets for various debugging methods. 1) Simple Feedback is a basic baseline that informs the LLM that the code is incorrect and asks it to fix the issue. 2) Self-Edit (Zhang et al., 2023) prompts the LLM to edit the code based on the execution results of the test cases. 3) Self-Debugging (Chen et al., 2023b) has two variants: Self-Debugging (Expl.) prompts the LLM to explain the generated code line-by-line, while Self-Debugging (Trace) asks the LLM to dry run the code for debugging. 4) LDB (Zhong et al., 2024) segments the code into basic blocks, functions or lines, and tracks variable values during runtime after each block to verify correctness against the task description. 5) Reflexion (Shinn et al., 2023) asks the LLM to reflect on the previous code given execution results and uses a memory buffer to enable iterative refinement.
@@
Implementation Details We generate seed programs for HumanEval and MBPP using the BigCode Evaluation Harness framework [^3] https://github.com/bigcode-project/bigcode-evaluation-harness. The specific versions of models used in our experiments are DeepSeek-Coder-V2-Lite-Instruct, CodeQwen1.5-7B-Chat, and Codestral-22B-v0.1. All experiments are conducted on NVIDIA A100 GPUs with 80GB memory. During debugging, we use the vLLM engine [^4] https://github.com/vllm-project/vllm to serve the LLMs, setting the maximum token length according to each LLM’s max length. Following Zhong et al. (2024), we limit the maximum number of debugging iterations to 10 for all methods on HumanEval, MBPP, and HumanEvalFix. To obtain visible test cases for HumanEval and HumanEvalFix, we extract the given visible test cases from the task description. For MBPP, we use the first test case of each problem as the visible test case and use the rest as hidden test cases, in line with the settings referenced from Chen et al. (2023b) and Zhong et al. (2024). Additionally, the sampling temperature that controls the randomness of LLM decoding is set to 0.8 in MGDebugger. For baseline methods that incorporate external feedback from the environments, we follow (Zhong et al., 2024) to execute the complete test code to collect runtime feedback.
@@ 3.3. Ablation Study (RQ2)
To understand the contribution of each component in MGDebugger and validate our design choices, we conduct an ablation study by systematically removing key components of our method: hierarchical code decomposition, LLM-simulated execution, and test case generation for subfunction debugging. Additionally, we explore the impact of replacing simulated execution with real execution traces collected with LDB (Zhong et al., 2024) to further validate our design choice. Each variant is evaluated on both the HumanEval and MBPP datasets using the DeepSeek-Coder-V2-Lite model.
@@
Furthermore, we observe that replacing simulated execution with real execution leads to a noticeable drop in performance, particularly in repair success rate (from 76.3% to 47.4% on HumanEval and from 39.0% to 36.0% on MBPP). This suggests that real execution traces, which include all detailed variable changes and execution steps, may introduce noise that distracts the model during debugging. In contrast, simulated execution allows MGDebugger to focus on key variables and express debugging steps in a more reasoning-oriented manner, leading to better code understanding and more accurate repairs. Furthermore, simulated execution offers greater flexibility for extension to complex programming environments where execution environments may not be easily accessible (Ni et al., 2024). This makes MGDebugger more adaptable to a wider range of scenarios.
@@ 3.4. Debugging Different Types of Bugs (RQ3)
Looking into details of different bug categories, MGDebugger shows a strong advantage in debugging bottom-level bugs, such as missing logic and excess logic. Missing logic refers to situations where essential code is omitted, preventing the solution from functioning correctly. Excess logic, on the other hand, involves unnecessary code that can lead to mistakes and confusion (Muennighoff et al., 2023). Other methods often struggle to identify and address these underlying issues because they treat the code holistically. This can lead to confusion over bottom-level details when dealing with complex logical errors. By contrast, the hierarchical decomposition in MGDebugger allows it to focus on different levels of code granularity. This enables more effective identification and correction of bugs. These results demonstrate the robustness and versatility of MGDebugger across various bug types.
@@ 3.5. Performance Consistency (RQ4)
Another important factor for LLM-based debugging is the number of debugging attempts. Iterative debugging retries allows LLMs to refine their corrections over multiple passes, potentially leading to better outcomes (Yin et al., 2024; Zhong et al., 2024). We aim to assess MGDebugger’s ability to improve over successive iterations. Following Zhong et al. (2024), we vary the number of debugging attempts from 1 to 10 using DeepSeek-Coder on different dataset and collect the cumulative repair success rate (RSR) scores for each method. We clarify that a single iteration refers to one complete bottom-up traversal of the function decomposition tree, including test case generation and simulated execution for all subfunctions. Even if a function is decomposed into multiple subfunctions, the process is counted as a single iteration if it involves one full pass from the leaves to the root.

### v4

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773255isbn: 979-8-4007-2025-3/26/04
@@ 1. Introduction
Large language models (LLMs) such as GPT-4 (openai2023gpt4), LLaMA (touvron2023llama), and DeepSeek-Coder (zhu2024deepseekcoderv2) have made significant advances in AI-assisted coding tasks (chen2021evaluating; li2022competitionlevel; just2014defects4j; shi2024between). Trained on vast corpora of text and code, LLMs can understand and generate code snippets for various programming tasks, ranging from simple data structures to complex algorithmic problems (li2022competitionlevel; wang2025dissect). These models have demonstrated proficiency in tasks such as code completion, bug detection, and even tackling competitive programming challenges (chen2021evaluating; zhang2024codeagent; wen2024learning).
While the code generated by large language models generally meets the requirements, it often contains critical errors that require human intervention to pass tests (liu2023your; dou2024what). This has gradually led to a new development paradigm: large models generate the code, while humans fix it. Therefore, the “last mile”, as well as the most crucial step, of code generation is how to efficiently repair the code generated by large models.
Numerous efforts have been made to debug LLM-generated code. The most popular way is to reuse the LLM generator to debug the generated code with the feedback from test case execution (chen2023teaching; zhong2024debug; hu2024leveraging). While these methods increase the pass rates, they treat the erroneous program as a holistic set of statements (i.e., as a single unit without considering the internal structure or modularity) (chen2023teaching; shinn2023reflexion; zhong2024debug; ding2024cycle) regardless of the varying types and levels of failures. Failures of test cases arise from different levels of factors, from low-level syntactic errors to high-level algorithmic flaws. A holistic treatment overlooks the internal structure of the code and limits the effectiveness of the debugging systems, especially when dealing with complex programs that need debugging across different modules (zeller2009why; tian2024debugbench).
@@
Extensive experiments across multiple models and benchmarks demonstrate that MGDebugger significantly outperforms existing debugging methods. Using open-source models like DeepSeek-Coder-V2-Lite, CodeQwen1.5, and Codestral, MGDebugger elevates accuracy from 75.6% to 94.5% on HumanEval (chen2021evaluating) and achieves a remarkable 97.6% repair success rate on HumanEvalFix (muennighoff2023octopack). The approach also generalizes well to state-of-the-art proprietary models, improving GPT-4o and Claude 3.5 Sonnet’s performance by 6.1% on HumanEval, reaching accuracies of 96.3% and 95.7% respectively.
@@
While MGDebugger is primarily designed to fix LLM-generated code and close the last mile of code generation, we also evaluate its generalizability to real-world software defects. When applied to the Defects4J dataset (just2014defects4j), MGDebugger successfully repairs 129 bugs in Defects4J v1.2 and 64 bugs in Defects4J v2.0, outperforming specialized program repair tools. This demonstrates that the hierarchical debugging principles with test case generation and LLM-simulated execution of MGDebugger extend effectively beyond LLM-generated code to human-written software with complex project structures.
@@ 2.2. Hierarchical Code Decomposition
Modularizing and decomposing complex code into smaller helper subfunctions has been proven to be helpful especially for large functions that are difficult to understand (jain2023llmassisted; zelikman2023parsel). To enable hierarchical debugging, we need to transform the input code into a tree-like structure of subfunctions.
Specifically, given an LLM-generated function $f$, we decompose it into a hierarchical structure of subfunctions denoted as $(f_{1},...,f_{n})$. We formally define a ”subfunction” as a contiguous sequence of statements $S=[s_{1},s_{2},...,s_{k}]$ within function $f$, such that $S$ represents a logically coherent and minimal unit of functionality that can be abstracted as a function $f_{i}(x)$ with well-defined inputs and outputs. These subfunctions can be organized as a tree $f_{\text{root}}=\text{TREE}(f_{\text{root}},\text{CHILD}(f_{\text{root}}))$, where $f_{\text{root}}$ represents the main function and $\text{CHILD}(f)$ denotes the set of subfunctions directly called by $f$. We leverage an LLM for the decomposition, adhering to three principles: (1) each subfunction represents the minimal reusable unit of code with a specific purpose, (2) higher-level functions call lower-level functions to achieve complex functionality, and (3) the overall structure facilitates isolated testing and debugging. As illustrated in Figure 1, the resulting tree-like structure allows us to isolate logical units of the code, enabling more focused debugging efforts across different levels of granularity (woodfield1981effect; isazadeh2017source). The prompt template for hierarchical decomposition is shown below:
@@ 2.3. Generating Test Cases for Subfunctions
Having obtained the hierarchy of subfunctions, we aim to verify the correctness of each subfunction. For this purpose, we generate test cases for each subfunction leveraging automatic unit test generation techniques (wang2021automatic; schafer2024empirical; liu2023rltf). For each subfunction $f_{i}\in f_{\text{root}}$, we generate a set of test cases $\mathcal{T}_{i}$. Following the problem settings from chen2023teaching and zhong2024debug, we assume that the public test cases for the main function $\mathcal{T}_{\text{pub}}$ have been provided, which is common in most code generation benchmarks (chen2021evaluating; hendrycks2021measuring; muennighoff2023octopack) [^2] Otherwise, we can use LLM-generated test cases instead.. We can leverage these test cases to derive a set of corresponding test cases for each subfunction.
@@ 2.4. Debugging Subfunctions with LLM-simulated Execution
One straightforward way to implement this process is to use an external Python executor to monitor runtime variable values (zhong2024debug). However, when debugging high-level functions, tracking variable values within lower-level subfunctions is often unnecessary, as their correctness is ensured by the bottom-up debugging methodology. Furthermore, directly collecting all execution traces from the external debugger can add unnecessary overhead and complexity to the process.
Inspired by the methodology in li2023chain, we propose an LLM-simulated code executor, which prompts LLM to act as a code interpreter and track the code execution. We choose LLM-simulated execution over real execution for several key reasons (la2024code; ni2024next; zheng2024opencodeinterpreter): (1) Many code snippets require specific environments, hardware, or dependencies that are difficult to set up or may not be available (e.g., embedded systems, special libraries), (2) Some code cannot be executed due to compilation errors or security concerns, necessitating sandboxing or virtual execution. As shown in Figure 2, we request the LLM to simulate the execution process, reasoning about key variables and their states at each step, and thoroughly analyzing the failed test cases. This eliminates the need for an external debugger, offering a more flexible and efficient debugging solution. In addition, the LLM can accurately identify where errors occur and grasp their surrounding context. The prompt template for debugging subfunction is shown below:
@@
12:  if pass ($\mathcal{R}_{f}$, $\mathcal{T}_{f}$) then
@@ 3.1. Setup
Models We select three state-of-the-art open-source LLMs ranging from 7B to 22B parameters as backbones for code generation and debugging in the experiments: CodeQwen1.5 (7B) (bai2023qwen), DeepSeek-Coder-V2-Lite (16B) (zhu2024deepseekcoderv2), and Codestral (22B) (mistralaiteam2024codestral). To validate the generalizability of MGDebugger, stronger general-prupose models like GPT-4o (openai2023gpt4), Claude 3.5 Sonnet (anthropic2024introducing), and LLaMA 3.1 70B (touvron2023llama) are also included in the Experiments.
@@
Datasets We conduct experiments on three widely used benchmarks for LLM-based code generation and debugging. HumanEval (chen2021evaluating) and MBPP (austin2021program) are two widely used benchmarks for evaluating code generation systems with 164 and 500 problems, respectively. The HumanEvalFix dataset (muennighoff2023octopack) consists of 164 buggy functions with six different bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse.
@@
Metrics We adopt two metrics to evaluate our method: 1) Accuracy (chen2023teaching; zhong2024debug), which measures the overall proportion of correct code samples among all generated code samples after debugging. A code is correct iff it passes all private test cases assigned to it. 2) Repair Success Rate (RSR) (yasunaga2021breakitfixit), which refers to the proportion of fixed code samples to the total number of buggy code samples. The former metric indicates to which extent MGDebugger can improve the code generation systems, and the latter metric directly reflects the effectiveness of MGDebugger in fixing bugs from LLM-generated code.
@@
Baselines We compare MGDebugger with eight state-of-the-art methods for debugging LLM-generated code. 0) No-Debugging baseline refers to directly using the LLM to generate code without applying any debugging method, following (chen2023teaching; zhong2024debug). This baseline creates the initial set of seed programs, the subset of buggy programs then become targets for various debugging methods. 1) Simple Feedback is a basic baseline that informs the LLM that the code is incorrect and asks it to fix the issue. 2) Self-Edit (zhang2023selfedit) prompts the LLM to edit the code based on the execution results of the test cases. 3) Self-Debugging (chen2023teaching) has two variants: Self-Debugging (Expl.) prompts the LLM to explain the generated code line-by-line, while Self-Debugging (Trace) asks the LLM to dry run the code for debugging. 4) LDB (zhong2024debug) segments the code into basic blocks, functions or lines, and tracks variable values during runtime after each block to verify correctness against the task description. 5) Reflexion (shinn2023reflexion) asks the LLM to reflect on the previous code given execution results and uses a memory buffer to enable iterative refinement.
@@
Implementation Details We generate seed programs for HumanEval and MBPP using the BigCode Evaluation Harness framework [^3] https://github.com/bigcode-project/bigcode-evaluation-harness. The specific versions of models used in our experiments are DeepSeek-Coder-V2-Lite-Instruct, CodeQwen1.5-7B-Chat, and Codestral-22B-v0.1. All experiments are conducted on NVIDIA A100 GPUs with 80GB memory. During debugging, we use the vLLM engine [^4] https://github.com/vllm-project/vllm to serve the LLMs, setting the maximum token length according to each LLM’s max length. Following zhong2024debug, we limit the maximum number of debugging iterations to 10 for all methods on HumanEval, MBPP, and HumanEvalFix. To obtain visible test cases for HumanEval and HumanEvalFix, we extract the given visible test cases from the task description. For MBPP, we use the first test case of each problem as the visible test case and use the rest as hidden test cases, in line with the settings referenced from chen2023teaching and zhong2024debug. Additionally, the sampling temperature that controls the randomness of LLM decoding is set to 0.8 in MGDebugger. For baseline methods that incorporate external feedback from the environments, we follow (zhong2024debug) to execute the complete test code to collect runtime feedback.
@@ 3.3. Ablation Study (RQ2)
To understand the contribution of each component in MGDebugger and validate our design choices, we conduct an ablation study by systematically removing key components of our method: hierarchical code decomposition, LLM-simulated execution, and test case generation for subfunction debugging. Additionally, we explore the impact of replacing simulated execution with real execution traces collected with LDB (zhong2024debug) to further validate our design choice. Each variant is evaluated on both the HumanEval and MBPP datasets using the DeepSeek-Coder-V2-Lite model.
@@
Furthermore, we observe that replacing simulated execution with real execution leads to a noticeable drop in performance, particularly in repair success rate (from 76.3% to 47.4% on HumanEval and from 39.0% to 36.0% on MBPP). This suggests that real execution traces, which include all detailed variable changes and execution steps, may introduce noise that distracts the model during debugging. In contrast, simulated execution allows MGDebugger to focus on key variables and express debugging steps in a more reasoning-oriented manner, leading to better code understanding and more accurate repairs. Furthermore, simulated execution offers greater flexibility for extension to complex programming environments where execution environments may not be easily accessible (ni2024next). This makes MGDebugger more adaptable to a wider range of scenarios.
@@ 3.4. Debugging Different Types of Bugs (RQ3)
Looking into details of different bug categories, MGDebugger shows a strong advantage in debugging bottom-level bugs, such as missing logic and excess logic. Missing logic refers to situations where essential code is omitted, preventing the solution from functioning correctly. Excess logic, on the other hand, involves unnecessary code that can lead to mistakes and confusion (muennighoff2023octopack). Other methods often struggle to identify and address these underlying issues because they treat the code holistically. This can lead to confusion over bottom-level details when dealing with complex logical errors. By contrast, the hierarchical decomposition in MGDebugger allows it to focus on different levels of code granularity. This enables more effective identification and correction of bugs. These results demonstrate the robustness and versatility of MGDebugger across various bug types.
@@ 3.5. Performance Consistency (RQ4)
Another important factor for LLM-based debugging is the number of debugging attempts. Iterative debugging retries allows LLMs to refine their corrections over multiple passes, potentially leading to better outcomes (yin2024thinkrepair; zhong2024debug). We aim to assess MGDebugger’s ability to improve over successive iterations. Following zhong2024debug, we vary the number of debugging attempts from 1 to 10 using DeepSeek-Coder on different dataset and collect the cumulative repair success rate (RSR) scores for each method. We clarify that a single iteration refers to one complete bottom-up traversal of the function decomposition tree, including test case generation and simulated execution for all subfunctions. Even if a function is decomposed into multiple subfunctions, the process is counted as a single iteration if it involves one full pass from the leaves to the root.